i. Suppose that the production function for compact disc player is

𝑄=100𝐿0.6𝐾0.4

Where 𝑄 is the total output, 𝐿 is the quantity of labor employed, and 𝐾 is the quantity of capital in place.

a) Calculate TP, AP, and MP for the sixth, seventh and eighth units of labour employed if capital is fixed at 240 units.


1
Expert's answer
2021-06-22T10:19:32-0400

Given 

"Q= 100L^{0.6}K^{0.4} ....... (1)"

Where Q is total output, L is the quantity of labor employed. K is quantity of capital. 

If capital is fixed at "K=240"

So 

"Q= 100L^{0.6} 240^{0.4}"

"Q= 895.54 L^{0.6} ...... (2)"


Average product of labor, 

Dividing output function by L

"AP(L) = 895.54 L^{(-0.4)} ....... (3)"

Marginal product of labor, differentiate equation 2 w r t L

"MP(L) =TP(n)-TP(n-1) .......... (4)"

 

When "L = 6"

"Q= 895.54{ (6^{0.6})}\\ 

Q= 2624.07"

So total product is"2624.07"

Average product 

"AP= 895.54 (6)^{(-0.4)}=437.345"

Marginal product of labor"MP(6) =Q(6)-Q(5) =2624.07-895.54 (5^{0.6}) =271.91"


When labor is L= 7 

Total product 

"Q= 895.54 (7^{0.6})\\ 

Q= 2878.35"

So total product is "2878.35"

Average product 

"AP= 895.54 (7)^{(-0.4)}=411.19"

Marginal product of labor

"MP(7) =TP(7)-TP(6) =254.28"

When L = 8

Total product 

"Q= 895.54 (8^{0.6}) \\

Q= 3118.45"

So total product is "3118.45"

Average product 

"AP= 895.54 (8)^{(-0.4)}=389.81"

Marginal product of labor

"MP(8) =TP(8)-TP(7)= 240.1"